Simplifiy and solve

3. 30 square root of -25 * 8 sqaure root of -49

4. 6 square root of -64 + 12 square root of -36

5. 4 sqaure root of -13 - 6 square root of -13
1
Expert's answer
2013-01-18T08:58:45-0500
3) 30 square root of -25 * 8 sqaure root of -49

30 * sqrt(-25) * 8 * sqrt(-49) =
= 30 * 5i * 8 *7i
= (30 * 5 * 8 *7) * (i*i)
= 8400 * (-1)
= -8400

-------------------------------------------------------
4) 6 square root of -64 + 12 square root of -36

6 * sqrt(-64) + 12 * sqrt(-36) =
= 6 * 8i + 12 *6i
= 48i + 72i
120i

-------------------------------------------------------
5) 4 square root of -13 - 6 square root of -13
4*sqrt(-13) - 6*sqrt(-13)
= -2 *sqrt(-13)
= -2 sqrt(13) *i